If f(x)=2x+5 , then the graphs of y=f(x) and y=squareroot f(x) will have two common points.

a. What are the y-coordinates of these common points?

b. Determine the x-coordinates of these common points, and then state the points.


1
Expert's answer
2021-05-13T18:10:55-0400

given

"f(x)=2x+5"

"let\space g(x)=\sqrt{\smash[b]{f(x)}}=\sqrt{\smash[b]{2x+5}}"

The graphs of "y=f(x)" and "y=g(x)" will intersect at points where "f(x)=g(x)"

"f(x)=g(x)\Rightarrow 2x+5=\sqrt{\smash[b]{2x+5}}"

"\Rightarrow2x+5-\sqrt{\smash[b]{2x+5}}=0"

"\Rightarrow\sqrt{\smash[b]{2x+5}}(\sqrt{\smash[b]{2x+5}}-1)=0"

"\Rightarrow\sqrt{\smash[b]{2x+5}}=0,\sqrt{\smash[b]{2x+5}}=1"

"\Rightarrow2x+5=0,2x+5=1"

"\Rightarrow x=-\frac{5}{2} \space ,\space x=-2"


The y-coordinates are

"x=-2\Rightarrow y=\sqrt{\smash[b]{2x+5}}=1"

"x=-\frac{5}{2}\Rightarrow y=\sqrt{\smash[b]{2x+5}}=0"

Hence, the x-coordinates of the points of intersection are "x=-\frac{5}{2},-2"

Hence, the x-coordinates of the points of intersection are "y=0,1"

The points of intersections are "(-2,1)" and "(-\frac{5}{2},0)"